OpenStudy (anonymous):

say that a and b are the lengths of a right triangle and c is the length of the hypotenuse. find b if a = 21 and c = 29.

OpenStudy (anonymous):

b=20

OpenStudy (smarty):

 So 21^2+B^2=29^2 441+B^2=841 Subtract 441 from both sies B^2=400 B=20

OpenStudy (anonymous):

\[c ^{2}=a ^{2}+b ^{2}\rightarrow 29^{2}=21^{2}+b ^{2}\rightarrow b ^{2}=29^{2}-21^{2}=400\rightarrow b =+\sqrt{400}=20\]